OpenStudy (anonymous):

Identify the vertex, focus, directrix, and min/max value of each. y+1=−(x−1)2

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you got the vertex? that is the easiest part

OpenStudy (anonymous):

yea i got 1,-1

OpenStudy (misty1212):

ok that is good

OpenStudy (misty1212):

you know this opens down right?

OpenStudy (anonymous):

down

OpenStudy (misty1212):

it looks like \(4p(y-k)=(x-h)^2\) so the vertex is \((1,-1\) like you said, and \(4p=-1\) so \(p=-\frac{1}{4}\) the focus is \(\frac{1}{4}\) units DOWN from \((1,-1)\) namely at \((1,-\frac{5}{4})\) and the directrix is \(\frac{1}{4}\) units up from the vertex, i.e at \(y=-\frac{3}{4}\)

OpenStudy (anonymous):

How do i get the max or min? would it be the max -2 and min 2?